OpenStudy (anonymous):

Given the parent function of f(x) = x3, what change will occur when the function is changed to f(x) - 3? Shift to the right 3 units Shift to the left 3 units Shift up 3 units Shift down 3 units @mathstudent55 Please help me!

OpenStudy (ageta):

f(x + 5) would shift the function to the LEFT 5 f(x-5) will shift it to the RIGHT 5

OpenStudy (ageta):

Shift to the left 3 units

OpenStudy (anonymous):

Think of it like all the y-values are being lowered by 3 so the shift is 3 units down

OpenStudy (anonymous):

thank you! @ageta

OpenStudy (anonymous):

The 3 is not part of the input of the function so it is not a left or right shift

OpenStudy (ageta):

u welcome

OpenStudy (anonymous):

No really. I'm serious. It is three units down

OpenStudy (anonymous):

Bob and Susie wash cars for extra money over the summer. Bob's income is determined by f(x) = 6x + 13, where x is the number of hours. Susie's income is g(x) = 4x + 18. If Bob and Susie were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Bob works 3 hours. Create the function h(x) and indicate if Bob will make more money working alone or by teaming with Susie. h(x) = 2x + 5, work alone h(x) = 2x + 5, team with Susie h(x) = 10x + 31, team with Susie h(x) = 10x + 31, work alone

OpenStudy (ageta):

h(x) = 10x + 31, team with Susie

OpenStudy (anonymous):

Solve 6 over x minus 6 equals the quantity of x over x minus 6, minus six halves for x and determine if the solution is extraneous or not. x = -6, extraneous x = -6, non-extraneous x = 6, non-extraneous x = 6, extraneous

OpenStudy (ageta):

first step would be making all the fractions have a common denominator To this multiply the first and second each with 2/2 Multiply the last one by (x-6)/(x-6)

OpenStudy (ageta):

would be 2x/2x-12

OpenStudy (ageta):

This is getting confusing.. 6 2. x. 2. 6. (x-6) ---- * ----- = ------ *. ------ - ---- * ----- (x-6). 2. (x-6). 2. 2. (x-6) Does this make sense? If so, can you simplify the numerator, because the denominator is all the same

OpenStudy (anonymous):

i got the answer thx :)

OpenStudy (ageta):

u will come